Deciding how many workers to hire: Assume that the initial price of shoes in this example...

  1. Deciding how many workers to hire: Assume that the initial price of shoes in this example is $30 per pair. What is the marginal revenue product for each worker? Fill in the following chart and graph each function..

No of workers

0

1

2

3

4

5

6

7

Total Output

Of shoes

0

9

19

27

33

36

38

37

Marginal Output of shoes

Marginal Revenue Product

  1. If it costs the firm $90 per worker per day, how many workers would be hired? Why?
  2. If the price of shoes went up to $45 per pair how many workers would be hired?
  3. If labor costs went up to $180/day and the price of shoes was $30 each, how many workers would be hired?
Answer #1
No. of workers 0 1 2 3 4 5 6 7
Total output of shoes 0 9 19 27 33 36 38 37
Marginal output of shoes -   9 -0 = 9 19 - 9 =10 27 - 19 = 8 6 3    2    -1
Marginal revenue product - 9*30 = 270 10*30 = 300 8*30 = 240 180 90 60 -30

a) Workers should be hired to the point where VMPL ( or Marginal revenue product)  = Labor cost per unit = $90.

This shows true when the no. of worker = 5

b) If price of shoes went upto $45

No. of workers 0 1   2 3 4 5 6 7
Marginal revenue product ( at $45) - 405 450 360 270 135 90 -45

Workers should be hired to the point where VMPL = Labor cost per unit = $90

This shows true when no. of workers = 6

c) Workers should be hired to the point where VMPL = Labor cost per unit = $180

This shows true when no. of workers = 4 (in thetable of part (a))